element 24 ground - and structer state for Configuration 1st this element is what about the elements 7,80 °F love, No, mg,
Answer #1

structure for a element is written as 2s+1LJ

where s=spin quantum number, it is either non-negative integer or zero only if number of electrons is even or half integer if number of electron is odd. spin multiplicityy=2S+1

L=total orbital angular momentum quantum number(a non-negative integer)

L= 0,1,2,3 respectively for S,P,D ,F and so on

J=Total angular momentum quantum number (Non -negative integer or half integer)=L+S,L+S-1,...........L-S

if orbital is less than half-filled then J=L-S

If more than half filled J=L+S

IF exactly halffilled or ful-filled then J=S

Structure is determined by valance orbital of element.

YFN = 15² 28² 283 hit L=1+0 1 = 0 so s S = 3/=3* Y2 C3 on paired 25 + 1 = 2x3 + 1 = 4 J: L+S 0 + 3/2 = 3/2 structure fore INDate:_//_Page nos yote = 13² 282 apo LLL L = 2x1 + 0x2+2x- = 2-2 = 0 so s S = = as t 1 = 1 J: Its : 0+0 EO Structeerce: As, YDate: +2002 = 152 252 2p6 352 3p6 43 11 10 TS S=0 25+ 1 = 2 x 0 + 1 = 1 JE 0 streecturce for this is o -